# Search for an eV-scale sterile neutrino with the first six detection units of KM3NeT/ORCA

### v1

@@ Abstract
The existence of an eV-scale sterile neutrino has been proposed to explain several anomalous experimental results obtained over the course of the past 25 years. The first search for such a sterile neutrino conducted with data from KM3NeT/ORCA — a water Cherenkov neutrino telescope under construction at the bottom of the Mediterranean Sea — is reported in this paper. GeV-scale atmospheric neutrino oscillations are measured by reconstructing the energy and arrival direction of up-going neutrinos that have traversed the Earth. This study is based on a data sample containing 5828 neutrino candidates collected with 6 detection units ($5\%$ of the complete detector), corresponding to an exposure of 433 kton-years. From the expected effect of an eV-scale sterile neutrino on the first $\nu_{\mu}\rightarrow\nu_{\tau}$ standard oscillation maximum, simultaneous constraints are put on the magnitude of the $U_{\mu 4}$ and $U_{\tau 4}$ mixing elements under the assumption $\Delta m^{2}_{41}=1$ eV2. The results are compatible with the absence of mixing between active neutrinos and a sterile state, with $|U_{\mu 4}|^{2}<0.138$ and $|U_{\tau 4}|^{2}<0.076$ at a $90\%$ confidence level. Such constraints are compatible with the results reported by other long-baseline experiments, and indicate that with KM3NeT/ORCA it is possible to bring crucial contributions to sterile neutrino searches in the coming years.
@@ 1 Introduction
Despite the success of the three-flavour PMNS paradigm at describing the neutrino oscillation data, several experimental results obtained at a relatively short baseline (SBL) remain largely unexplained aceroWhitePaperLight2023 . All these anomalies, despite originating from experiments with different neutrino sources, detection technologies, oscillation channels, baselines and neutrino energies, can be individually interpreted as oscillations driven by a squared-mass difference $\Delta m^{2}\sim\mathcal{O}(1)$ eV2. This value of $\Delta m^{2}$ is several orders of magnitude larger than the measured values of $\Delta m^{2}_{31}$ and $\Delta m^{2}_{21}$. Thus, to explain the SBL anomalies in terms of oscillations, a simple extension of the standard model is to assume the existence of a fourth, eV-scale, neutrino state. However, cosmological constraints exist on the effective number of relativistic neutrino species lesgourguesMassiveNeutrinosCosmology2006 ; navasReviewParticlePhysics2024 . In addition, the LEP experiments obtained compelling evidence that only three flavours of light neutrinos that couple to the weak interaction — known as active neutrinos — exist PrecisionElectroweakMeasurements2006 . [^1] In that context, light means with a mass lower than half of the mass of the Z boson. Thus, an additional eV-scale neutrino would have to be insensitive to the weak interaction, or sterile. In the following, this model will be referred to as the 3++1 model.
@@ 5.3 Statistical methods
The test statistic used to build the confidence interval of the parameters of interest around their best-fit values is the logarithm of the likelihood ratio between each point of the phase-space of interest $\vec{x}$ and the best-fit point $\vec{\widehat{x}}$, written $2\Delta\ln\mathcal{L}$ in the following. The confidence intervals of the parameters of interest are built in a frequentist way. Using Wilks’ theorem wilksLargeSampleDistributionLikelihood1938 , the allowed region of the parameter space given a certain confidence level is obtained by comparing the value of $2\Delta\ln\mathcal{L}$ with the corresponding quantile of a $\chi^{2}$ distribution with two degrees of freedom, as two is the dimension of the parameter space of interest.
@@ 5.4 Fitting procedure
For each point of the phase space of interest, equation 9 is minimised using 2 starting points for 4 nuisance parameters, to avoid getting stuck in local minima: normal and inverted ordering ($\Delta m^{2}_{31}=\{2.541\times 10^{-3},-2.496\times 10^{-3}\}~\mathrm{eV^{2}}$); $\theta_{23}$ lower and upper octant ($\theta_{23}=\{40^{\circ},50^{\circ}\}$); energy scale $E_{s}$ below and above 1.0 ($E_{s}=\{0.95,1.05\}$); and $\delta_{24}$ below and above $180^{\circ}$ ($\delta_{24}=\{90^{\circ},270^{\circ}\}$). This produces a total of 16 distinct starting points. The two start values for $\Delta m_{31}^{2}$, $\theta_{23}$ and $E_{s}$ are common to the standard oscillation analysis of ORCA6 aielloMeasurementNeutrinoOscillation2024 . For $\delta_{24}$, the use of two start values is motivated by the observation of local minima when computing the $2\Delta\ln\mathcal{L}$ profile of that parameter. In each fit, the parameter space is restricted to the NMO, $\theta_{23}$ octant and $(E_{s}-1)$ sign corresponding to the start value. No limit is put on $\delta_{24}$.
@@ 6 Results
The ORCA6 data sample is analysed under the $\Delta m_{41}^{2}=1$ eV2 hypothesis, although the results also hold for $\Delta m_{41}^{2}>1$ eV2 for which faster oscillations are unobservable. Simultaneous constraints on the mixing of the $\nu_{\mu}$ and $\nu_{\tau}$ states with the sterile state are obtained by scanning the $(|U_{\mu 4}|^{2},|U_{\tau 4}|^{2})$ space, with $U_{e4}=0$ (i.e. $\theta_{14}=\delta_{14}=0$). In that case, the magnitude of the $U_{\mu 4}$ and $U_{\tau 4}$ mixing elements is expressed as:
@@ 6.1 Best-fit point
Figure 5: Muon neutrino survival probability as a function of the neutrino energy under various scenarios for $\cos\theta=-1.0$. The three-flavour standard oscillation model assuming the nominal value for $\theta_{23}=49.1\degree$ (dashed black) is compared with the three-flavour model with $\theta_{23}=59\degree$ (dashed grey) and with the 3++1 models with $\sin^{2}\theta_{24}=0.1$ (pink) and $\sin^{2}\theta_{34}=0.15$ (yellow), for $\Delta m_{41}^{2}=1$ eV2. All other sterile parameters are set to 0.
@@ 6.3 Comparison with existing measurements
The ORCA6 upper limits at $90\%$ CL are compared with the results of other experiments in figure 6. They are competitive with these existing results, especially concerning the limit on the $\nu_{\tau}$ mixing with the sterile state: only DeepCore, using $7.5$ years of data, puts a stronger constraint on $|U_{\tau 4}|^{2}$. For $|U_{\mu 4}|^{2}$, ANTARES, IceCube, DeepCore and Super-Kamiokande all put stronger constraints than ORCA6, but these results were obtained with several years of data-taking from complete detectors, while the ORCA6 measurement was obtained with 1.4 years of data recorded with a detector that was $5\%$ of its final size. All analyses considered here treat $\delta_{24}$ as a free nuisance parameter, with the exception of Super-Kamiokande and IceCube. The Super-Kamiokande analysis fixes $\delta_{24}=0$, leading to more restrictive limits than with a free $\delta_{24}$. The IceCube analysis relies on the TeV disappearance feature which is less sensitive to $\delta_{24}$. It uses $\delta_{24}=\pi$, as this choice was found to yield the most conservative limits. The performance of KM3NeT/ORCA is due to its good sensitivity to $\nu_{\mu}$ in the 203020-30 GeV region which allows for the monitoring of the first $\nu_{\mu}\rightarrow\nu_{\tau}$ standard oscillation maximum. Other detectors have to rely on weaker effects at lower energy (e.g. Super-Kamiokande) or are optimised for slightly higher energies (e.g. ANTARES).
Figure 6: Comparison of the upper limits at $90\%$ CL on $|U_{\mu 4}|^{2}$ and $|U_{\tau 4}|^{2}$ obtained by ORCA6 with previous measurements from Super-Kamiokande super-kamiokandecollaborationLimitsSterileNeutrino2015 , NOvA thenovacollaborationSearchActiveSterileAntineutrino2021 , ANTARES albertMeasuringAtmosphericNeutrino2019 , IceCube abbasiExplorationMassSplitting2024a and DeepCore icecubecollaborationSearchLightSterile2024 . The excluded regions are on the top right side of the contours.
@@ 7 Conclusion
The first search for an eV-scale sterile neutrino performed with the KM3NeT/ORCA detector is presented in this article. In this analysis, a data sample collected with 6 detection units ($5\%$ of the complete detector size) is used, corresponding to an exposure of 433 kton-years and containing 5828 neutrino candidates. Constraints on the mixing of a hypothetical eV-scale sterile neutrino with the $\nu_{\mu}$ and $\nu_{\tau}$ states are established under the assumption $\Delta m^{2}_{41}=1$ eV2. The results are compatible with the three-flavour standard model, and the upper limits obtained on $|U_{\mu 4}|^{2}$ and $|U_{\tau 4}|^{2}$ are compatible with the results reported by other experiments.

### v2

@@ Abstract
The existence of an eV-scale sterile neutrino has been proposed to explain several anomalous experimental results obtained over the course of the past 25 years. The first search for such a sterile neutrino conducted with data from KM3NeT/ORCA — a water Cherenkov neutrino telescope under construction at the bottom of the Mediterranean Sea — is reported in this paper. GeV-scale atmospheric neutrino oscillations are measured by reconstructing the energy and arrival direction of up-going neutrinos that have traversed the Earth. This study is based on a data sample containing 5828 neutrino candidates collected with 6 detection units ($5\%$ of the complete detector), corresponding to an exposure of 433 kton-years. From the expected effect of an eV-scale sterile neutrino on the first $\nu_{\mu}\rightarrow\nu_{\tau}$ standard oscillation maximum, simultaneous constraints are put on the magnitude of the $U_{\mu 4}$ and $U_{\tau 4}$ mixing elements assuming $\Delta m^{2}_{41}\geq 1$ eV2. The results are compatible with the absence of mixing between active neutrinos and a sterile state, with $|U_{\mu 4}|^{2}<0.138$ and $|U_{\tau 4}|^{2}<0.076$ at a $90\%$ confidence level. Such constraints are compatible with the results reported by other long-baseline experiments, and indicate that with KM3NeT/ORCA it is possible to bring crucial contributions to sterile neutrino searches in the coming years.
@@ 1 Introduction
Despite the success of the three-flavour PMNS paradigm at describing the neutrino oscillation data, several experimental results obtained at a relatively short baseline (SBL) remain largely unexplained aceroWhitePaperLight2023. All these anomalies, despite originating from experiments with different neutrino sources, detection technologies, oscillation channels, baselines and neutrino energies, can be individually interpreted as oscillations driven by a squared-mass difference $\Delta m^{2}\sim\mathcal{O}(1)$ eV2. This value of $\Delta m^{2}$ is several orders of magnitude larger than the measured values of $\Delta m^{2}_{31}$ and $\Delta m^{2}_{21}$. Thus, to explain the SBL anomalies in terms of oscillations, a simple extension of the standard model is to assume the existence of a fourth, eV-scale, neutrino state. However, cosmological constraints exist on the effective number of relativistic neutrino species lesgourguesMassiveNeutrinosCosmology2006; particledatagroupReviewParticlePhysics2024. In addition, the LEP experiments obtained compelling evidence that only three flavours of light neutrinos that couple to the weak interaction — known as active neutrinos — exist PrecisionElectroweakMeasurements2006.111In that context, light means with a mass lower than half of the mass of the Z boson. Thus, an additional eV-scale neutrino would have to be insensitive to the weak interaction, or sterile. In the following, this model will be referred to as the 3++1 model.
@@ 5.3 Statistical methods
The test statistic used to build the confidence interval of the parameters of interest around their best-fit values is the logarithm of the likelihood ratio between each point of the phase-space of interest $\vec{x}$ and the best-fit point $\vec{\widehat{x}}$, written $2\Delta\ln\mathcal{L}$ in the following. The confidence intervals of the parameters of interest are built in a frequentist way. Using Wilks’ theorem wilksLargeSampleDistributionLikelihood1938, and not a full Feldman-Cousins confidence interval derivation feldmanUnifiedApproachClassical1998 due to computing constraints, the allowed region of the parameter space given a certain confidence level is obtained by comparing the value of $2\Delta\ln\mathcal{L}$ with the corresponding quantile of a $\chi^{2}$ distribution with two degrees of freedom, as two is the dimension of the parameter space of interest.
@@ 5.4 Fitting procedure
For each point of the phase space of interest, equation 9 is minimised using two starting points for four nuisance parameters, to avoid getting stuck in local minima: normal and inverted ordering ($\Delta m^{2}_{31}=\{2.541\times 10^{-3},-2.496\times 10^{-3}\}~\mathrm{eV^{2}}$); $\theta_{23}$ lower and upper octant ($\theta_{23}=\{40^{\circ},50^{\circ}\}$); energy scale $E_{s}$ below and above 1.0 ($E_{s}=\{0.95,1.05\}$); and $\delta_{24}$ below and above $180^{\circ}$ ($\delta_{24}=\{90^{\circ},270^{\circ}\}$). This produces a total of 16 distinct starting points. The two start values for $\Delta m_{31}^{2}$, $\theta_{23}$ and $E_{s}$ are common to the standard oscillation analysis of ORCA6 aielloMeasurementNeutrinoOscillation2024. For $\delta_{24}$, the use of two start values is motivated by the observation of local minima when computing the $2\Delta\ln\mathcal{L}$ profile of that parameter. In each fit, the parameter space is restricted to the NMO, $\theta_{23}$ octant and $(E_{s}-1)$ sign corresponding to the start value. No limit is put on $\delta_{24}$.
@@ 6 Results
The ORCA6 data sample is analysed under the $\Delta m_{41}^{2}=1$ eV2 hypothesis; the results hold for $\Delta m_{41}^{2}>1$ eV2 for which faster oscillations are unobservable. Simultaneous constraints on the mixing of the $\nu_{\mu}$ and $\nu_{\tau}$ states with the sterile state are obtained by scanning the $(|U_{\mu 4}|^{2},|U_{\tau 4}|^{2})$ space, with $U_{e4}=0$ (i.e. $\theta_{14}=\delta_{14}=0$). In that case, the magnitude of the $U_{\mu 4}$ and $U_{\tau 4}$ mixing elements is expressed as:
@@ 6.1 Best-fit point
Figure 5: Muon neutrino survival probability as a function of the neutrino energy under various scenarios for $\cos\theta=-1.0$. The three-flavour standard oscillation model assuming the nominal value for $\theta_{23}=49.1\degree$ (dashed black) is compared with the three-flavour model with $\theta_{23}=59\degree$ (dotted black) and with the 3++1 models with $\sin^{2}\theta_{24}=0.1$ (blue) and $\sin^{2}\theta_{34}=0.15$ (orange), for $\Delta m_{41}^{2}=1$ eV2. All other sterile parameters are set to 0.
@@ 6.3 Comparison with existing measurements
The ORCA6 upper limits at $90\%$ CL are compared with the results of other experiments in figure 6. Table 3 summarizes how each analysis treats the 3+1 and standard oscillation parameters relevant for this measurement. All consider eV-scale sterile neutrinos, but do not report exactly the same $\Delta m_{41}^{2}$ validity range. The IceCube result is obtained by profiling over $\Delta m_{41}^{2}$ between 0.1 eV2 and 50 eV2, with a best fit found at 5.0 eV2. All analyses considered here treat $\delta_{24}$ as a free nuisance parameter, with the exception of Super-Kamiokande and IceCube. The Super-Kamiokande analysis fixes $\delta_{24}=0$, leading to more restrictive limits than with a free $\delta_{24}$. The IceCube analysis relies on the TeV disappearance feature which is less sensitive to $\delta_{24}$. It uses $\delta_{24}=\pi$, as this choice was found to yield the most conservative limits. Regarding the mass ordering, only the ORCA6 and the ANTARES analyses consider both orderings. Finally, due to the partial degeneracies between $\theta_{23}$ and the $\theta_{24}$ and $\theta_{34}$ mixing angles (see figure 5), analyses lead with no constraint on this parameter (as done in this work, and for the DeepCore and ANTARES analysis) yield more conservative limits than the other analyses.
Experiment $\Delta m_{41}^{2}$ range (eV2) $\delta_{24}$ Mass ordering $\Delta m_{31}^{2}$ $\theta_{23}$
This work $\geq 1.0$ Free Both Prior from dayabaycollaborationPrecisionMeasurementReactor2023 Free
DeepCore icecubecollaborationSearchLightSterile2024 $\geq 1.0$ Free NO only Free
IceCube abbasiExplorationMassSplitting2024a Profiled in $[0.1,50]$ $\pi$ Unknown Unknown
NOvA thenovacollaborationSearchActiveSterileAntineutrino2021 $[0.05,0.5]$ Free NO only Fixed from particledatagroupReviewParticlePhysics2018
ANTARES albertMeasuringAtmosphericNeutrino2019 $\geq 0.5$ Free Both Free
SK super-kamiokandecollaborationLimitsSterileNeutrino2015 $\geq 0.1$ =0=0 NO only Prior from t2kcollaborationPreciseMeasurementNeutrino2014
Table 3: Summary of the configurations used for the analyses compared in figure 6. The assumed $\Delta m_{41}^{2}$ validity range, and the treatment of $\delta_{24}$, the mass ordering, $\Delta m_{31}^{2}$ and $\theta_{23}$ in the fits are reported. Other oscillation parameters do not have a significant impact on the $|U_{\mu 4}|^{2}$ and $|U_{\tau 4}|^{2}$ measurement.
The ORCA6 limits reported in figure 6 are competitive with the existing results, especially concerning the limit on the $\nu_{\tau}$ mixing with the sterile state: only DeepCore, using $7.5$ years of data, puts a stronger constraint on $|U_{\tau 4}|^{2}$. For $|U_{\mu 4}|^{2}$, ANTARES, IceCube, DeepCore and Super-Kamiokande all put stronger constraints than ORCA6, but these results were obtained with several years of data-taking from complete detectors, while the ORCA6 measurement was obtained with 1.4 years of data recorded with a detector that was $5\%$ of its final size. The performance of KM3NeT/ORCA is due to its good sensitivity to $\nu_{\mu}$ in the 203020-30 GeV region which allows for the monitoring of the first $\nu_{\mu}\rightarrow\nu_{\tau}$ standard oscillation maximum. Other detectors have to rely on weaker effects at lower energy (e.g. Super-Kamiokande) or are optimised for slightly higher energies (e.g. ANTARES).
Figure 6: Comparison of the upper limits at $90\%$ CL on $|U_{\mu 4}|^{2}$ and $|U_{\tau 4}|^{2}$ obtained by ORCA6 with previous measurements from Super-Kamiokande (SK) super-kamiokandecollaborationLimitsSterileNeutrino2015, NOvA (Nv) thenovacollaborationSearchActiveSterileAntineutrino2021, ANTARES (AN) albertMeasuringAtmosphericNeutrino2019, IceCube (IC) abbasiExplorationMassSplitting2024a and DeepCore (DC) icecubecollaborationSearchLightSterile2024. The excluded regions are on the top right side of the contours.
@@ 7 Conclusion
The first search for an eV-scale sterile neutrino performed with the KM3NeT/ORCA detector is presented in this article. In this analysis, a data sample collected with 6 detection units ($5\%$ of the complete detector size) is used, corresponding to an exposure of 433 kton-years and containing 5828 neutrino candidates. Constraints on the mixing of a hypothetical eV-scale sterile neutrino with the $\nu_{\mu}$ and $\nu_{\tau}$ states are established for $\Delta m^{2}_{41}\geq 1$ eV2. The results are compatible with the three-flavour standard model, and the upper limits obtained on $|U_{\mu 4}|^{2}$ and $|U_{\tau 4}|^{2}$ are compatible with the results reported by other experiments.